# High-Fidelity And Complex Test Data Generation For Real-World SQL Code Generation Services

### v2

@@ III-C Prompt builder
The test data generation prompt is constructed using the following template in Figure 5. The template contains some system instructions to specify the task and also to prevent common hallucination problems, like repeating the same column more than once, carefully validating and matching the column/field names to the schema.
@@
1. Generate the proto column name only if present in the proto file. 2. If you have already generated data for a field, please move to the next field. 3. Generate non-zero values for every numeric field. ”””
system_instruction = f”””
{system_prompt_prefix}
{data_generation_signals}
user_instruction = f”””
Task: Given the following proto description, generate
{number_of_data_points} rows of sample data for col_names columns.
{user_input}
Proto/schema Description:
{proto_description}
mock_data_gen_prompt =
f”””
system:{system_instruction}
user:{user_instruction}
model:
1---------------------------------------------------- 2# Without column grouping: 3 4json[{ 5 "contract_created_date": "2023-08-15", 6 "contract_created_month": "2023-08" 7 "contract_created_year": 2023, 8 "contract_effective_date": "2023-04-01", 9 "contract_closed_date": "2023-03-30", 10 "contract_executed_date": "2023-03-29", 11 "contract_expiry_date": "2024-03-31", 12 "fully_signed_time": "2023-03-29T10:00:00Z", "contract_termination_date": "2023-12-31", 13}, 14...] 15---------------------------------------------------- 16# With column grouping: 17 18json[{ 19 "contract_created_date": "2023-10-27", 20 "contract_created_month": 10, 21 "contract_created_year": 2023, 22 "contract_effective_date": "2023-11-01", 23 "contract_closed_date": "2024-10-31", 24 "contract_executed_date": "2023-10-30", 25 "contract_expiry_date": "2025-10-31", 26 "fully_signed_time": "2023-10-30T10:00:00Z", 27 "contract_termination_date": "2024-05-15", 28 "contract_termination_date_option": "SpecificDate", 29 "contract_months_from_effective_date": 12, 30 "simba_contract_create_date": "2023-10-27", 31}, 32...]
@@
1---------------------------------------------------- 2# column annotation 3 4"terminated_by_clm_system" // "Ifthecontractwasterminated,thisfieldindicatestheContractLifecycleManagementSystemofthecontract(s)responsibleforitstermination.Thisfieldispartofthesourceinformationaboutthiscontract’slineagewithintheContractFamilyLinkage." 5 6---------------------------------------------------- 7# Generated column value without annotation: 8"terminated_by_clm_system": "TerminationSystem", 9 10---------------------------------------------------- 11# Generated column value with annotation: 12"terminated_by_clm_system": "InternalCLMSystemCode",
@@ III-E Data validator
Instructions for validation: Check the values required for a column under the constraints section. Check for the same column in the data section. If the column is not present, please return NOT PRESENT. If the columns is present, please check if the data is present in the constraint for that column. If point 4 is satisfied, please print VALID. If point 4 is not satisfied, please print NOT VALID. user_prompt = f”””
Task: Given the following Data and Constraints, validate the data and return one of VALID, NOT VALID, NOT PRESENT.
Data:
{data}
Constraints:
{constraints}
Semantic constraints and validation is not easy to capture with conventional approaches, and we use the LLM for semantic validation (LLM-as-a-judge [33]) and we pass the SQL constraints extracted from the pre-processor module along with the generated test data, as part of the template in Figure 8. Table III summarizes various syntactic and semantic data validation rules enforced by the generator; the generator retries the generation up to three times along with the error message upon any validation failure, including the semantic ones. In section IV-C, we evaluate how often the LLM hallucinates and returns semantically invalid data that passes the validation checks.
@@ IV Evaluation results
In this section, we present evaluation results of our proposed LLM-based test data generator. The evaluation investigates three key aspects that highlight the efficacy and advantages of our approach. First, we assess the test coverage achieved by the generated data, focusing on its ability to facilitate more complex SQL query testing. Second, we evaluate the completeness and correctness of the generation for complex nested data structures – as commonly encountered in our SQL code generation workloads. This evaluation is based on the criteria used in Table 2, ensuring a quantitative assessment of the generated data’s structural integrity and fidelity to the defined schemas and protos. Through these evaluations, we demonstrate the superiority of our approach in creating high-quality, comprehensive test datasets. Lastly, we evaluate the effectiveness of our data generation methodology in comparison to production data sampling. Based on a limited set of authorized production tables from restricted domains, we perform a comparative analysis of the generated data’s ability to provide meaningful test data with respect to the target test SQL query/function, covering all the edge cases for semantically relevant predicate columns.
@@ IV-B Test coverage with more complex SQL query testing
We assess the test coverage achieved by the generated test data. Our generation method can generate test data for complex schema with deeply nested structures. This facilitates more comprehensive testing, enabling more complex SQL test cases that were disabled due to the lack of such complex test data. Table IV presents improvement (%) in test coverage and test accuracy – we report the improvement instead of the raw test coverage values. The improvement in NL2SQL testing coverage is more significant at $69.59\%$ because many available test cases were disabled due to restricted production data access. Notice that the generated test data also improves NL2SQL test accuracy. To our surprise, there were many cases where manually curated production test data that failed to differentiate predicted and golden SQL queries. For instance, some aggregate queries with wrong predicted filter predicates resulted in the correct answers (false positives); many test queries with INNER JOIN were flagged as incorrect (false negatives) against the manually curated test data, whereas adding sufficient test data for all the edge cases for the relevant columns used in the SQL query (Section VI-B) resulted in correct matching results. In SQL Code Assistant there are a lot more disabled test cases (mostly disabled because it is challenging to generate complex nested test data instances manually), and we expect to roll out new generated test data for those in the near future, which should increase the coverage further.
TABLE IV: Test coverage and test accuracy changes (%) enabled by complex data generation and with more complex SQL test cases. NL2SQL test coverage improvement is much significant because many complex test cases were disabled due to restricted production table access before; NL2SQL test accuracy also improved significantly, since the manually curated test data allowed many false positives and false negatives.
@@ IV-C Generation quality evaluation with complex nested data structures
Next, we evaluate the completeness and correctness of the generation for complex nested data structures. Specifically, we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types.
@@
Figure 9 presents our structural integrity & adherence to complex data schema evaluation results. We use a subset of key evaluation criteria derived from the ones used in Table II – we excluded the criteria for primary key generation, which is always true for our test data. The evaluation scores are normalized (0 to 1) and the higher the better. Overall, our generated test data adheres very well to complex schema with deeply nested column types encountered in production SQL code generation services. There are some cases where nested field (proto) definitions are stale or removed, in which case the LLM guesses the field names and types for the inaccessible nested proto structures. Note that the scores are much higher (and near perfect) without such invalid proto definition cases w/o invalid proto. Note that the correct nesting levels scores for both with and without invalid proto are the same because we skip over those invalid nested structures in the aggregation – since we cannot assess the correct nesting levels for unknown proto. We typically see 3-level of nesting in complex SQL testing scenarios, and as deep as 5 levels.
@@ IV-D Semantic data constraints validation with LLM-as-a-judge
The generator extracts semantic data constraints, like which column values are relevant to the given SQL query in the pre-processing step. We use the LLM as LLM-as-a-judge to validate the generated test data against the extracted semantic constraints. If some generated data instances fail to satisfy the constraints according to the LLM, then we retry with an additional hint that the previous generation failed to satisfy the failed semantic constraints. In all test cases, we have successfully generated test data instances after 1 or 2 retries. Interestingly, we observed that the LLM can hallucinate and actually pass invalid data instances as valid. Table V illustrates how generating multiple data instances per test case can limit the hallucination error rates, defined as the proportion of test cases with no semantically valid test data instances after $n$ instances generated. We report the hallucination error rates for NL2SQL test cases only because the unit test cases for SQL Code Assistant are too simple and contains simple semantic constraints, like COUNTIF(experiment_group = “GROUP A”) AS group_a_count. While most semantic constraint violations are resolved with a single additional data generation attempt ($n=2$), a few constraints remain challenging for the LLM-as-a-judge to validate. Specifically, complex datetime filter constraints that require type casting, time zone conversion, and date extraction consistently failed even after five attempts ($n=5$). For example, consider the constraint DATE(TIMESTAMP_SECONDS(CAST(status.time_processed _sec AS INT64)), ’America/Los_Angeles’) $\geq$ “2023-07-01”. Given a generated value like ”1678886400” for time_processed_sec, the validator correctly performed the initial conversion, recognizing that the timestamp corresponds to the date “2023-03-15”. However, it then failed to perform the final step of evaluating whether this date satisfied the constraint (i.e., that “2023-03-15” is not greater than or equal to “2023-07-01”). A potential remedy involves invoking a reasoning model or chain-of-thought prompting [34], though this significantly increases generation time. We plan to explore more cost-efficient ways to address these hallucination errors in future work.
TABLE V: LLM hallucination error rates for semantic data validation with repeated data generation ($n$, the number of test data instances) per test case. By generating multiple instances $n>1$, we get at least one test instance that satisfies the semantic constraints for most of the test cases. Semantic validation for SQL Code Assistant test cases was perfect, since the unit test cases did not involve any complex semantic constraints.
@@
Here we evaluate the effectiveness of our data generation methodology in comparison to production data sampling. We use one production table authorized for testing to generate data samples of different sizes. To compare the quality of our generated test data against the production samples, we compare the number of distinct values of the relevant predicates/columns extracted from the question and golden SQL query. We use a test case submitted by an anonymous developer, “How many Shorts videos about Gaming were uploaded by (identified as) female creators in the United Kingdom(GB) during Q1 2024?,” for the test production table. Note that evaluating the NL2SQL output result is not meaningful as it should return different answers depending on different samples. There are two relevant columns also used as predicates, describing different country codes and semantic tags for the Shorts videos.
@@
Figure 10 presents the number of distinct values for the (a) country codes column and (b) semantic tags column. Because the question is asking for the (identified as) female case, we also track the counts by (identified as) male and (identified as) female. In case of country codes, the production samples Prod include a more diverse set of countries than the Mock ones. This is OK because the question is looking for the answer in the United Kingdom (GB) specifically – and it is likely that our generation is less diverse because of this. At the same time, we also notice an imbalanced sampling between Prod/M and Prod/F in the production samples, which is not desirable as the question seeks to answer for the (identified as) female case. In case of semantic tags, which can entail different genres and keywords for different gaming contents (FPS, MMO, RPG, tutorial, review, etc.), our generated test data Mock does include a more diverse set of values. In both cases, the data is well balanced between the two select genders. Note that the test data is generated with syntax and semantic validations as described in Section III-E, but the values can still be unrealistic as it is not learned from the actual production data (not available). Our generation approach also takes the test query/function into account to generate data that is more semantically aligned to the query. We discuss another example in Section VI-B.
@@ V Related works
Test data generation for complex SQL testing: Test data generation is crucial for SQL query validation, query rewriting, automated query grading, and NL2SQL tasks. Common strategies involve using manually curated test data, production data sampling, constraints based data generation for table schema and SQL query [5, 1, 3, 4, 6]. Production data sampling can provide the most realistic test data for any complex schema, but it is often prohibited or infeasible in industrial contexts where service providers lack access to user data due to stringent privacy and security protections/policies. One of the key challenges in generating realistic high-fidelity test data is handling more complex nested schema structures and their inter-dependencies. Many SQL equivalence testing and constraints based SQL test data generation do not support arbitrarily nested structures and nested fields (array, struct and/or proto types) [11, 5] – limiting the test use cases to less complex SQL queries.
Model-based tabular data generation: Despite the emerging interest in LLMs for test data generation [16, 19, 20, 21], existing approaches often fail to fully address the complex testing requirements of industrial applications. While LLMs can provide realistic textual data and synthesize tabular data that adheres to schema format [27], the out-of-the-box LLMs fall short at capturing subtle nuances and relationships betwen data atrributes [25]. We also found that the vanilla LLM generation becomes more challenging and incomplete for a large nested schema (Table II). Other deep generative model-based techniques have been used for synthetic tabular data [29, 30, 31]. While producing realistic data examples that closely approximate the underlying distributions, the traditional tabular data generation solutions focus on flat schema without nested fields (struct, array, proto). Another challenge is to source good training datasets for learning the tabular generation model, since the production data is often restricted in the real-world testing environments. In this work, we propose a complex test data generation method based on a pre-trained LLM, gemini, and without any post-training.

### v3

@@ III-C Prompt builder
The test data generation prompt is constructed using the following template in Figure 5. The template contains some system instructions to specify the task and also to prevent common hallucination problems, like repeating the same column more than once, carefully validating and matching the column/field names to the schema. While structural hallucination (e.g., incorrect column/field names) are easier to detect and address almost perfectly in the subsequent post-processing and data validation steps, semantic hallucination (e.g., generating test data in fiscal year 2023, while the NL2SQL test question is about fiscal year 2025) is much harder to address completely, as further discussed in Section III-E.
@@
1. Generate the proto column name only if present in the proto file. 2. If you have already generated data for a field, please move to the next field. 3. Generate non-zero values for every numeric field. ‘‘‘ system_instruction=f""" {system_prompt_prefix} {data_generation_signals} """ user_instruction=f""" Task: Given the following proto description, generate {number_of_data_points} rows of sample data for {col_names} columns. {user_input} Proto/schema Description: {proto_description} """ mock_data_gen_prompt=f""" system:{system_instruction} user:{user_instruction} model: ‘‘‘
// Without column grouping: ---------------------------------------------------- json[{ "contract_created_date": "2023-08-15", "contract_created_month": "2023-08" "contract_created_year": 2023, "contract_effective_date": "2023-04-01", "contract_closed_date": "2023-03-30", "contract_executed_date": "2023-03-29", "contract_expiry_date": "2024-03-31", "fully_signed_time": "2023-03-29T10:00:00Z", "contract_termination_date": "2023-12-31", }, ...] // With column grouping: ---------------------------------------------------- json[{ "contract_created_date": "2023-10-27", "contract_created_month": 10, "contract_created_year": 2023, "contract_effective_date": "2023-11-01", "contract_closed_date": "2024-10-31", "contract_executed_date": "2023-10-30", "contract_expiry_date": "2025-10-31", "fully_signed_time": "2023-10-30T10:00:00Z", "contract_termination_date": "2024-05-15", "contract_termination_date_option": "SpecificDate", "contract_months_from_effective_date": 12, "simba_contract_create_date": "2023-10-27", }, ...]
@@
// Column annotation ---------------------------------------------------- "terminated_by_clm_system" // "If the contract was terminated, this field indicates the Contract Lifecycle Management System of the contract(s) responsible for its termination. This field is part of the source information about this contract’s lineage within the Contract Family Linkage." // Generated column value without annotation: ---------------------------------------------------- "terminated_by_clm_system": "TerminationSystem", ---------------------------------------------------- # Generated column value with annotation: "terminated_by_clm_system": "InternalCLMSystemCode",
@@ III-E Data validator
Instructions for validation: Check the values required for a column under the constraints section. Check for the same column in the data section. If the column is not present, please return NOT PRESENT. If the columns is present, please check if the data is present in the constraint for that column. If point 4 is satisfied, please print VALID. If point 4 is not satisfied, please print NOT VALID. ‘‘‘ user_prompt = f""" Task: Given the following Data and Constraints, validate the data and return one of VALID, NOT VALID, NOT PRESENT. Data: {data} Constraints: {constraints} """ ‘‘‘
Semantic constraints and validation is not easy to capture with conventional approaches, and we use the LLM for semantic validation (LLM-as-a-judge [31]) and we pass the SQL constraints extracted from the pre-processor module along with the generated test data, as part of the template in Figure 8. Table III summarizes various syntactic and semantic data validation rules enforced by the generator; the generator retries the generation up to three times along with the error message upon any validation failure, including the semantic ones. In section IV-C, we evaluate how often the LLM hallucinates and returns semantically invalid data that passes the validation checks (Table VI).
@@ IV Evaluation results
We evaluate our LLM-based test data generator by investigating three key aspects that demonstrate its advantage in creating high-quality, comprehensive test datasets:
  • Complex Query Support: We assess the test coverage achieved, focusing on the data’s utility for testing complex SQL queries.
  • Schema Adherence: We quantitatively evaluate the data’s structural integrity and fidelity to the complex nested schemas often encountered in our SQL code generation workloads, using criteria defined in Table II.
  • Practical Effectiveness: We compare our generated data against limited authorized production samples to demonstrate its effectiveness in covering all semantically relevant edge cases for target test SQL queries/functions.
@@ IV-B Test coverage with more complex SQL query testing
We assess the test coverage achieved by the generated test data. Our generation method can generate test data for complex schema with deeply nested structures. This facilitates more comprehensive testing, enabling more complex SQL test cases that were disabled due to the lack of such complex test data.
Table IV presents improvement (%) in test coverage and test accuracy – we report the improvement instead of the raw test coverage values. The improvement in NL2SQL testing coverage is more significant at $69.59\%$ because many available test cases were disabled due to restricted production data access. Notice that the generated test data also improves NL2SQL test accuracy. To our surprise, there were many cases where manually curated production test data that failed to differentiate predicted and golden SQL queries. For instance, some aggregate queries with wrong predicted filter predicates resulted in the correct answers (false positives); many test queries with INNER JOIN were flagged as incorrect (false negatives) against the manually curated test data, whereas adding sufficient test data for all the edge cases for the relevant columns used in the SQL query (Section VI-B) resulted in correct results. In SQL Code Assistant there are a lot more disabled test cases due to lack of test data– it is challenging to generate complex nested test data instances manually. We expect to roll out new generated test data for those in the near future, which should increase the coverage further.
Table V presents the ablation study on NL2SQL test data generation by removing each component from our data generation pipeline. Column grouping and data validation steps are more relevant for semantic validity and the test coverage by test executability remains unchanged without them. We do not include SQL Code Assistant, since the unit test cases consists of simple selection and projection queries – SQL parsing and schema fetching are sufficient to populate the relevant schema structures and fields to get the full coverage improvement.
TABLE IV: Test coverage and test accuracy changes (%) enabled by complex data generation and with more complex SQL test cases.
@@
TABLE V: Relative test coverage difference without each component for test data generation in NL2SQL.
Component Removed Coverage Difference (%)
SQL parsing 100.00-100.00
Schema fetching 100.00-100.00
Proto description 47.76-47.76
Column grouping $0.00$
Post-processing 20.15-20.15
Data Validation $0.00$
@@ IV-C Generation quality evaluation with complex nested data structures
Next, we evaluate the completeness and correctness of the generation for complex nested data formats. Specifically, we examine the generated test data quality in terms of its structural integrity to the complex schema with columns with deeply nested field types. Figure 9 shows the normalized results (0 to 1, higher is better), using key criteria from Table II – we excluded the criteria for primary key generation, which is always true for our test data. Despite issues with some inaccessible schemas in production environment, our data generation successfully produces test data compliant with complex, real-world nested structures.
Overall, our generated test data adheres very well to complex schema with deeply nested column types encountered in production SQL code generation services. There are some cases where nested field (proto) definitions are stale or removed, in which case the LLM guesses the field names and types for the inaccessible nested proto structures. Note that the scores are much higher (and near perfect) without such invalid proto definition cases w/o invalid proto. Note that the correct nesting levels scores for both with and without invalid proto are the same because we skip over those invalid nested structures in the aggregation – since we cannot assess the correct nesting levels for unknown proto. We typically see 3-level of nesting in complex SQL testing scenarios, and as deep as 5 levels.
@@ IV-D Semantic data constraints validation with LLM-as-a-judge
The generator extracts semantic data constraints, like which column values are relevant to the given SQL query in the pre-processing step. We use the LLM as LLM-as-a-judge to validate the generated test data against the extracted semantic constraints. If some generated data instances fail to satisfy the constraints according to the LLM, then we retry with an additional hint that the previous generation failed to satisfy the failed semantic constraints. In all test cases, we have successfully generated test data instances after 1 or 2 retries. Interestingly, we observed that the LLM can hallucinate and actually pass invalid data instances as valid. Table VI illustrates how generating multiple data instances per test case can limit the hallucination error rates, defined as the proportion of test cases with no semantically valid test data instances after $n$ instances generated. We report the hallucination error rates for NL2SQL test cases only because the unit test cases for SQL Code Assistant are too simple and contains simple semantic constraints, like COUNTIF(experiment_group = “GROUP A”) AS group_a_count. It is important to note that without schema-level annotations (e.g., column grouping), the pipeline often generates semantically irrelevant test data instances (more than 88% of the time). Interestingly, the few successful test cases were simple aggregation or select queries that required only a handful of syntactically valid rows in the tables. While most semantic constraint violations are resolved with a single additional data generation attempt ($n=2$), a few constraints remain challenging for the LLM-as-a-judge to validate. Specifically, complex datetime filter constraints that require type casting, time zone conversion, and date extraction consistently failed even after five attempts ($n=5$). For example, consider the constraint DATE(TIMESTAMP_SECONDS(CAST(status.time_processed _sec AS INT64)), ’America/Los_Angeles’) $\geq$ “2023-07-01”. Given a generated value like ”1678886400” for time_processed_sec, the validator correctly performed the initial conversion, recognizing that the timestamp corresponds to the date “2023-03-15”. However, it then failed to perform the final step of evaluating whether this date satisfied the constraint (i.e., that “2023-03-15” is not greater than or equal to “2023-07-01”). A potential remedy involves invoking a reasoning model or chain-of-thought prompting [32], though this significantly increases generation time. We plan to explore more cost-efficient ways to address these hallucination errors in future work.
TABLE VI: LLM hallucination error rates for semantic data validation with repeated data generation ($n$, the number of test data instances) per test case. By generating multiple instances $n>1$, we get at least one test instance that satisfies the semantic constraints for most of the test cases. Semantic validation for SQL Code Assistant test cases was perfect, since the unit test cases did not involve any complex semantic constraints.
@@
Here we evaluate the effectiveness of our data generation methodology in comparison to production data sampling. We use one production table authorized for testing to generate data samples of different sizes. To compare the quality of our generated test data against the production samples, we compare the number of distinct values of the relevant predicates/columns extracted from the question and golden SQL query. We use a test case submitted by an anonymous developer, “How many Shorts videos about Gaming were uploaded by (identified as) female creators in the United Kingdom(GB) during Q1 2024?,” for the test production table. There are two relevant columns also used as predicates, describing different country codes and semantic tags for the Shorts videos.
@@
Figure 10 presents the number of distinct values for the (a) country codes column and (b) semantic tags column. Because the question is asking for the (identified as) female case, we also track the counts by (identified as) male and (identified as) female. In case of country codes, the production samples Prod include a more diverse set of countries than the Mock ones. This is OK because the question is looking for the answer in the United Kingdom (GB) specifically – and it is likely that our generation is less diverse because of this. At the same time, we also notice an imbalanced sampling between Prod/M and Prod/F in the production samples, which is not desirable as the question seeks to answer for the (identified as) female case. In case of semantic tags, which can entail different genres and keywords for different gaming contents (FPS, MMO, RPG, tutorial, review, etc.), our generated test data Mock does include a more diverse set of values. In both cases, the data is well balanced between the two select genders. Note that the test data is generated with syntax and semantic validations as described in Section III-E, but the values can still be irrelevant as it is not learned from the actual production data (not available). Our generation approach also takes the test query/function into account to generate data that is more semantically aligned to the query. We discuss another example in Section VI-B.
@@ V Related works
Test data generation for complex SQL testing: Test data generation is crucial for SQL query validation, query rewriting, automated query grading, and NL2SQL tasks. Common strategies involve using manually curated test data, production data sampling, constraints based data generation for table schema and SQL query [5, 1, 3, 4, 6]. Recent research [20] has explored the utility of LLMs for generating artifacts critical to database testing and development, like diverse and complex SQL test queries. This area, along with the broader application of LLMs in test case generation [33, 34], is attracting significant interest. This line of research, and the related benchmarks, is primarily concerned with leveraging LLMs to generate a diverse test SQL queries beyond a handful of simple SQL equivalent pairs; however, this is currently done using rather simple, non-production database schema.
In contrast, our work addresses a distinct but equally critical challenge: the complex and relevant generation of the test data for large complex, nested production schema structures – with relevant data instances that adhere to intricate real-world constraints. This is essential for effectively unit-testing complex SQL functions and evaluating the robustness of NL2SQL systems at Google. The key challenges in generating high-fidelity test data is handling more complex nested schema structures and their inter-dependencies. Many SQL equivalence testing and constraints based SQL test data generation do not support arbitrarily nested structures and nested fields (array, struct and/or proto types) [9, 5] – limiting the test use cases to less complex SQL queries. Production data sampling can provide the most realistic high-fidelity test data for any complex schema, but it is often prohibited or infeasible in industrial contexts where service providers lack access to user data due to stringent privacy and security protections/policies.
Model-based tabular data generation: Despite emerging interest in using LLMs for test data generation [14, 23, 24, 25], existing approaches often fail to meet the complex requirements of industrial testing. While LLMs can handle basic schema adherence [21], out-of-the-box models struggle with subtle data relationships [18] and become incomplete for large, nested schemas (Table II).
Similarly, traditional deep generative models [29, 30, 35] focus on flat schemas and require extensive training data, which is often restricted in real-world environments.
To address these limitations, we propose a complex test data generation method based on the pre-trained LLM, gemini, requiring no post-training.